Find the missing probability.
P(A∩B)=3/100,P(B|A)=3/20,P(A)=?

Answer by Edwin McCravy(20055)

P(A∩B)=3/100 = 0.03 = P(B∩A),

P(B|A)=3/20 = 0.15 = P(B∩A)/P(A) = 0.03/P(A)

0.15 = 0.03/P(A)

0.15P(A) = 0.03

P(A) =0.03/0.15 = 0.2 = 2/10 = 1/5

Edwin